OpenStudy (wasiqss):

Show that {A + A (transpose) } is hermatian and {A- A transpose) is Skew hermatian

OpenStudy (wasiqss):

I HAVE no idea how to approach this

OpenStudy (anonymous):

do you know the proof that \(\frac{f(x)+f(-x)}{2}\) is even and \(\frac{f(x)-f(-x)}{2}\) is odd? because if i remember correctly this is amost idential. you work from the definition of hermitian and skew

OpenStudy (wasiqss):

no im very bad at these stuff can you plz show the proof of my quesion

OpenStudy (experimentx):

G = (A + A') G' = (A + A')' = A' + A ...

OpenStudy (wasiqss):

But @experiment what to write in t a prrof im asked to submit the full proof

OpenStudy (anonymous):

This only holds iff A is Hermatian. For a proof, just take the transpose and use the definition of Hermatian and skew Hermatian matrices.

OpenStudy (anonymous):

And @satellite73 I think the proof you are reminiscing about is showing that every square matrix can be written as the sum of a symmetric and skew symmetric matrix.

OpenStudy (anonymous):

yes you are right, sorry

OpenStudy (anonymous):

oops, This works for any square matrix. My apologies!.

OpenStudy (wasiqss):

so ffm what is the right answer?

OpenStudy (anonymous):

The right answer is the proof that sat was reminiscing also and should works here!

OpenStudy (wasiqss):

so i should write the proof of sattelite?

OpenStudy (anonymous):

Also the decomposition is unique.

OpenStudy (anonymous):

No, that was just a hint.

OpenStudy (wasiqss):

sorry bit confuses so what should i write ,whose solution?

OpenStudy (anonymous):

Sorry, I had to sleep at that time. Btw do you mean symmetric?

OpenStudy (anonymous):

otherwise the result doesn't hold.. or do you mean conjugate transpose?